calculate the lateral area of a rectangular prism with a base with measure of 5.6in and 4.3in and a height of 10 inches?
Answer by fcabanski(1391)

The lateral area is the area of only the vertical sides. The formula for lateral area of a rectangular prism is Ph where P is the perimeter of the base so P = 2L + 2W.


In this case P = 2(5.6) + 2(4.3) = 11.2+8.6 = 19.8. The h = 10.


Lateral area is 19.8 * 10 = 198 square inches.

Hope the solution helped. Sometimes you need more than a solution. Contact fcabanski@hotmail.com for online, private tutoring, or personalized problem solving (quick for groups of problems.)